# Understanding Flow Behaviors of Supercooled Liquids by Embodying Solid-Liquid Duality at Particle Level

### v1

@@ V.1 PEL perspective on $\tau_{\mathrm{LC}}$ and $\tau_{\mathrm{sh}}$
Figure 24 (a) shows $f(t)$ at different $\dot{\gamma}$. To calculate $f(t)$, we set $E_{\mathrm{b}}$ in Eq. V.1 with the most probable value of the equilibrium distribution of $E_{\mathrm{b}}$ of sample A, denoted as $E_{\mathrm{b}}^{\prime}$. As $\dot{\gamma}$ increases, $f(t)$ as a whole shifts to smaller $t$. Moreover, the form of $f(t)$ also changes with $\dot{\gamma}$. In Fig. 24 (b), we plot the scaled $f(t)$, denoted as $f^{*}(t^{*})$, where $t^{*}=t/\bar{t}(E_{\mathrm{b}}^{\prime},\dot{\gamma})$ and $f^{*}=f/f_{\mathrm{max}}$ ($f_{\mathrm{max}}$ is the maximum value of $f(t)$). It is seen that $f(t)$ gradually changes from an exponential form to a gaussian-like form as $\dot{\gamma}$ increases. Particularly, for intermediate shear rates, $f(t)$ exhibits a bimodal form composed of an exponential component and a gaussian component, with the gaussian component being enhanced by $\dot{\gamma}$.
Figure 24: Functional forms of $f(t)$ as various $\dot{\gamma}$. (a) $f(t)$ in linear scale. (b) $f^{*}(t^{*})$, where $t^{*}=t/\bar{t}(E_{\mathrm{b}}^{\prime},\dot{\gamma})$ and $f^{*}=f/f_{\mathrm{max}}$. (c) $f(t)$ in log-log scale. The form at equilibrium $f_{\mathrm{eq}}(t)$ is highlighted by a thick line. For each $\dot{\gamma}$, we mark the position $t_{\mathrm{c}}$ at which $f(t)$ deviates from $f_{\mathrm{eq}}(t)$ by a circle with its color the same to that of the corresponding $\dot{\gamma}$. (d) The proportion of slow particles found from the PEL analysis $p_{\mathrm{slow}}^{\prime}$ as a function of $\dot{\gamma}$. Inset: $t_{\mathrm{c}}$ as a function of $\dot{\gamma}$ in log-log scale (symbols). The dashed line denotes the relation $t_{\mathrm{c}}\sim\dot{\gamma}^{-1}$.
@@
The alternation of the form of $f(t)$ corresponds to the observation in Secs. II and III that more particles are accelerated by shear as $\dot{\gamma}$ increases. To clarify this point, we plot $f(t)$ in the log-log scale in Fig. 24 (c). The one at equilibrium, $f_{\mathrm{eq}}(t)$, is highlighted by a thick solid line. For $f(t)$ with finite $\dot{\gamma}$, all of them align with $f_{\mathrm{eq}}(t)$ at small $t$, and deviate from $f_{\mathrm{eq}}(t)$ at large $t$. We denote the point at which $f(t)$ deviates from $f_{\mathrm{eq}}(t)$ as $t_{\mathrm{c}}$, and mark them in Fig. 24 (c). Evidently, $t_{\mathrm{c}}$ can be viewed as the boundary between fast and slow groups: For $t<t_{\mathrm{c}}$, particles are not accelerated by external shear and should be classified as fast; while for $t>t_{\mathrm{c}}$, particles are accelerated and should be classified as slow. Then, we calculate the percentage of slow particles by $p_{\mathrm{slow}}^{\prime}=\int_{t_{\mathrm{c}}}^{\infty}f(t)\,\mathrm{d}t$, and show the result in Fig. 24 (d). The S-shaped dependence of $p_{\mathrm{slow}}^{\prime}$ on $\lg{\dot{\gamma}}$ is similar to the results of $p_{\mathrm{slow}}$ shown in Fig. 13, which are directly found from the MD results. The inset of Fig. 24 (d) shows $t_{\mathrm{c}}$ as a function of $\dot{\gamma}$ in the log-log scale, which reveals a clear relation of $t_{\mathrm{c}}\propto\dot{\gamma}^{-1}$. This behavior is also consistent with the MD results shown in Fig. 3 (c) and (d).
@@ V.2 Steady shear
In the preceding subsection, we model $\tau_{\mathrm{LC}}$, $\tau_{\mathrm{sh}}$ and their relation as the interplay among local configuration, convection and thermal activation through the viewpoint of PEL. The discussion is primarily based on the start-up deformation represented by the behavior with respect to a single basin. In this subsection, we will further leverage the PEL picture to rationalize our findings on the steady shear.
@@
where $P=P(E_{\mathrm{b}},\gamma,t)$ is the probability that a coherent subsystem experiences a local energy barrier of $E_{\mathrm{b}}$ and a local strain of $\gamma$ at a given time $t$. In our picture, $P(E_{\mathrm{b}},\gamma,t)$ describes the dynamics at the particle level. The first term on the right-hand side of Eq. 19 represents the convective effect. The second term represents the activation from the distorted basin, where the energy barrier is given by $E_{\mathrm{b}}-E(\gamma)$. Here, the form of $E(\gamma)$ and the attempt frequency $\Gamma$ are set to those used in the preceding subsection. The third term represents the process that the yielded subsystems rearrange to new local equilibrium configurations. $\Gamma(t)$ is the total yielding rate, which is given by:
@@
$\rho(E_{\mathrm{b}})$ is the density of states for new barriers. Here, it is assumed that $\rho(E_{\mathrm{b}})$ does not depend on $\dot{\gamma}$ [23]. Therefore, $\rho(E_{\mathrm{b}})$ can be obtained from the dynamics of equilibrium state as follows. First, we measure the distribution of $\tau_{\mathrm{LC}}$ at equilibrium $H_{\mathrm{eq}}(\tau_{\mathrm{LC}})$. Second, assuming the activation picture, we combine $H_{\mathrm{eq}}(\tau_{\mathrm{LC}})$ and Eq. 14 to find the distribution of barrier height $P_{\mathrm{b}}(E_{\mathrm{b}})$. Then, $\rho(E_{\mathrm{b}})$ can be found by:
@@
The model given by Eqs. 1921 provides a framework for quantifying the flow behaviors under shear.
As demonstrated in Sec. IV, the dynamics and rheology of supercooled liquids are strongly correlated to the local structure. It is interesting to understand this correlation from the PEL perspective. We recall that local structural terms $g_{0,i}^{0}$ and $g_{2,i}^{-2}$ are most relevant to local dynamics compared to terms with other combinations of $l$ and $m$. In the PEL picture, local dynamics is determined by the local energy barrier $E_{\mathrm{b}}$ and the local strain $\gamma$. As for $g_{0,i}^{0}$, it represents the local packing surrounding particle $i$. Notably, a recent work [26] has established the quantitative connection between the energy barrier for particles and the local packing efficiency characterized by the particle-wise radial particle density, which is similar to $g_{0,i}^{0}$, for the equilibrium state. Following this result, it is clear that the importance of $g_{0,i}^{0}$ for local dynamics comes from its remarkable relation to the local energy barrier $E_{\mathrm{b}}$. Moreover, we find that the terms with $l=0$ and $m=0$ alter with $\dot{\gamma}$ only slightly. This result might justify the above approximation that $\rho(E_{\mathrm{b}})$ is $\dot{\gamma}$-independent.
Our emphasis will be on local structural terms with $l=2$ and $m=-2$. This kind of terms has received less attention in previous studies of the equilibrium state. However, they are very important for shear states as shown in Sec. IV. To understand this observation, we first refer to the conclusion that the strain of an affine shear $\gamma_{\mathrm{aff}}$ is proportional to $g_{2}^{-2}(r)$ [140, 141]:
@@
where $g_{2}^{-2}(r)$ is the ($l=2,m=-2$) coefficient of the spherical harmonic expansion of the global pair distribution function $g(\bm{r})$, and $g_{\mathrm{eq}}(r)$ is $g(\bm{r})$ at the undeformed state. Equation 22 suggests that $g_{2,i}^{-2}$ directly reflects the strain of the local configuration of particle $i$. Now, it is crucial to verify that if the strain variable $\gamma$ in the PEL framework, which can be calculated with the SGR model, can describe the configurational distortion reflected by $g_{2,i}^{-2}$ under steady shear. For this purpose, we perform the following calculations by employing the SGR model given by Eqs. 1921.
Figure 25: (a) and (b) show the results of the structure feature $Y$, defined by Eq. 26, as a function of the distance $r$ at various $\dot{\gamma}$ for sample A and sample B, respectively. The first seven peaks of $Y(r)$ are marked. (c) and (d) show the magnitudes of the first seven peaks of $Y(r)$ (symbols) at different $\dot{\gamma}$ for sample A and sample B, respectively. The results of $\bar{\gamma}$ (dashed lines), defined by Eq. 24, at different $\dot{\gamma}$ are also plotted. In (c) and (d), the values of peak magnitudes of $Y(r)$ are scaled with a numerical factor to compare with $\bar{\gamma}$.
According to Eq. 19, it can be demonstrated that the distribution of local strain $\gamma$ under steady shear is given by the following equation [23, 139]:
@@
In Sec. IV.1, we find that the second generation of $g_{l,i}^{m}$ well connects the local structure to local dynamics. Here, we adopt the second generation of $g_{2,i}^{-2}$ to represent the local distortion. Moreover, for the sake of simplicity, we do not differentiate particle species when calculating $g_{2,i}^{-2}$ with Eq. 9. With these considerations, we construct the following vector for describing the distortion of the local configuration of particle $i$:
@@
We plot the results of $Y(r_{k})$ with a series of $k$ values at different $\dot{\gamma}$ for sample A and sample B in Fig. 25 (a) and (b), respectively. $Y(r_{k})$ displays an oscillation as $r_{k}$ increases, which is similar to the $r$-dependence of $g_{2}^{-2}(r)$ [42]. As expected, the amplitude of $Y$ grows with $\dot{\gamma}$. Figure 25 (c) and (d) display the magnitudes of the first seven peaks of $Y(r_{k})$, which are marked in Fig. 25 (a) and (b), as a function of $\dot{\gamma}$ for sample A and sample B, respectively. Meanwhile, we also plot the results of $\bar{\gamma}$ found from Eq. 24 for both samples in Fig. 25 (c) and (d). It is seen that the magnitudes of the peaks of $Y(r_{k})$ and $\bar{\gamma}$ exhibit highly consistent dependences on $\dot{\gamma}$. This consistency holds for different peaks of $Y(r_{k})$ within $r_{k}<3.5$. At larger $r_{k}$, this consistency weakens, implying the localization of such deformation coherency. The remarkable proportionality between $Y(r_{k})$ and $\bar{\gamma}$ within $r_{k}<3.5$ implies a nice correspondence between $g_{2,i}^{-2}$ and the local strain $\gamma$ in the PEL framework.
In Sec. IV.1, we show that local structural parameters constructed by $g_{0,i}^{0}$ and $g_{2,i}^{-2}$, which respectively reflect the energy barrier height $E_{\mathrm{b}}$ and the strain $\gamma$ of the particle-level basin in the PEL framework according to the above discussion, well predict $\tau_{\mathrm{LC}}$ for all shear rates through a simple linear regression method. To explore the intrinsic mechanism of this correlation between local structure and dynamics, we will evaluate the dependence of $\tau_{\mathrm{LC}}$ on $\dot{\gamma}$ with the SGR framework. Under steady shear, the distribution of local strain $\gamma$ and barrier height $E_{\mathrm{b}}$ in the SGR model is given by:
@@
Figure 26 gives $\bar{\tau}_{\mathrm{LC}}$ as a function of $\dot{\gamma}$ for both samples. We also plot the average $\bar{\tau}_{\mathrm{LC}}$ measured directly from MD data in Fig. 26. It is seen that the results found by the SGR model are well consistent with the MD results.
@@
Figure 26: (a) and (b) show $\bar{\tau}_{\text{LC}}$ (dashed lines), defined by Eq. 29, as a function of $\dot{\gamma}$ for sample A and sample B, respectively. Meanwhile, the average $\tau_{\mathrm{LC}}$ measured from the MD data (symbols) are also shown. For MD results, squares denote the results at the equilibrium state, and circles denote the results for finite shear rates.
@@
The results shown in Figs. 25 and 26 suggest a clear picture from the PEL perspective for the universal correlation between local structure and $\tau_{\mathrm{LC}}$ at all shear rates shown in Sec. IV.1. The local structural terms related to $g_{0,i}^{0}$ and $g_{2,i}^{-2}$ respectively associate with the local energy barrier height $E_{\mathrm{b}}$ and the local strain $\gamma$ of the particle-level basin in the PEL picture. $\tau_{\mathrm{LC}}$ is the waiting time to jump out from the basin characterized by $\gamma$ and $E_{\mathrm{b}}$. The convective effect is reflected by the tilt of the basin, as illustrated in Fig. 21. Under steady shear, the collective effects of thermal activation and convection are described by the SGR model. With these considerations, we can unify the values of $\tau_{\mathrm{LC}}$ from configurations at different shear rates.
@@
Figure 27: (a) and (b) show $P_{\mathrm{ac}}(\gamma_{\mathrm{ac}})$, given by Eq. 30, at various $\dot{\gamma}$ for sample A and sample B, respectively. Here, the values of $P_{\mathrm{ac}}$ are scaled as $P_{\mathrm{ac}}^{*}(\gamma_{\mathrm{ac}})=P_{\mathrm{ac}}(\gamma_{\mathrm{ac}})/P_{\mathrm{ac,max}}(\gamma_{\mathrm{ac}})$. The vertical dashed line in each panel denotes the effective range of $P_{\mathrm{ac}}(\gamma_{\mathrm{ac}})$ at the lowest $\dot{\gamma}$. (c) and (d) show the results of the percentage of shear-driven particles $p_{\mathrm{sd}}$ (dashed lines), defined by Eq. 31, at different $\dot{\gamma}$ for sample A and sample B, respectively. The results of the proportion of slow particles $p_{\mathrm{slow}}$, which has been shown in Fig. 13, are replotted (symbols).
@@
In our framework, the solid-liquid duality, which is based on the grouping of slow/fast particles, is the key to connecting $\tau_{\mathrm{LC}}$ to rheological behaviors. Especially, this grouping is also strongly correlated to the local structure featured by terms associated with $g_{0,i}^{0}$ and $g_{2,i}^{-2}$, as demonstrated in Sec. IV.2. Thus, it is important to examine if the PEL picture can give such grouping for conditions of steady shear. To explore this problem, we introduce the concept of activation strain $\gamma_{\mathrm{ac}}$, which represents the strain at which particles jump out from the basin. Particular attention is paid to the distribution of $\gamma_{\mathrm{ac}}$, denoted as $P_{\mathrm{ac}}(\gamma_{\mathrm{ac}})$. Under steady shear, $P_{\mathrm{ac}}(\gamma_{\mathrm{ac}})$ is given by:
@@
Figure 27 (a) and (b) give the $P_{\mathrm{ac}}(\gamma_{\mathrm{ac}})$ of sample A and sample B, respectively. For the sake of clarity, we plot the scaled results $P_{\mathrm{ac}}^{*}(\gamma_{\mathrm{ac}})=P_{\mathrm{ac}}(\gamma_{\mathrm{ac}})/P_{\mathrm{ac,max}}(\gamma_{\mathrm{ac}})$, where $P_{\mathrm{ac,max}}(\gamma_{\mathrm{ac}})$ is the maximum value of $P_{\mathrm{ac}}(\gamma_{\mathrm{ac}})$. At the lowest $\dot{\gamma}$ shown here (in the Newtonian regime), $P_{\mathrm{ac}}(\gamma_{\mathrm{ac}})$ is exponential-like. In Fig. 27 (a) and (b), we use vertical dashed lines to mark the effective range of $P_{\mathrm{ac}}(\gamma_{\mathrm{ac}})$ at the lowest $\dot{\gamma}$, denoted as $\gamma_{\mathrm{ac,c}}$. As $\dot{\gamma}$ increases, $P_{\mathrm{ac}}(\gamma_{\mathrm{ac}})$ shifts to the regime of larger $\gamma_{\mathrm{ac}}$. More importantly, $P_{\mathrm{ac}}(\gamma_{\mathrm{ac}})$ gradually exhibits a bimodal feature composed of an exponential-like component at small $\gamma_{\mathrm{ac}}$ regime and a gaussian-like component at large $\gamma_{\mathrm{ac}}$ regime, with the latter being enhanced by shear. The emergence of the gaussian-like component at large $\gamma_{\mathrm{ac}}$ manifests the change of the way of relaxation. For particles corresponding to the gaussian-like component, their local configurations are significantly strained by external shear before relaxation. Qualitatively, these particles are likely slow particles: Their dynamics are slower than external shear, so that there is enough time for their local configurations to be strongly strained by external shear. On the contrary, remaining particles can be considered as likely fast particles, since there is no enough time for external shear to drive their local configurations to larger strains.
As seen in Fig. 27 (a) and (b), $\gamma_{\mathrm{ac,c}}$ gives a rough boundary between the exponential-like component and the gaussian-like component. Thus, we can evaluate the percentage of the gaussian-like component as follows:

### v2

@@ V.1 Shear-facilitated-activation picture of $\tau_{\mathrm{LC}}$ and $\tau_{\mathrm{sh}}$
Figure 24 (a) shows $f(t)$ at different $\dot{\gamma}$. To calculate $f(t)$, we set $E_{\mathrm{b}}$ in equation V.1 with the most probable value of the equilibrium distribution of $E_{\mathrm{b}}$ of sample A, denoted as $E_{\mathrm{b}}^{\prime}$. As $\dot{\gamma}$ increases, $f(t)$ as a whole shifts to smaller $t$. Moreover, the form of $f(t)$ also changes with $\dot{\gamma}$. In figure 24 (b), we plot the scaled $f(t)$, denoted as $f^{*}(t^{*})$, where $t^{*}=t/\bar{t}(E_{\mathrm{b}}^{\prime},\dot{\gamma})$ and $f^{*}=f/f_{\mathrm{max}}$ ($f_{\mathrm{max}}$ is the maximum value of $f(t)$). It is seen that $f(t)$ gradually changes from an exponential form to a gaussian-like form as $\dot{\gamma}$ increases. Particularly, for intermediate shear rates, $f(t)$ exhibits a bimodal form composed of an exponential component and a gaussian component, with the gaussian component being enhanced by $\dot{\gamma}$.
@@
The alternation of the form of $f(t)$ corresponds to the observation in sections II and III that more particles are facilitated by shear as $\dot{\gamma}$ increases. To clarify this point, we plot $f(t)$ in the log-log scale in figure 24 (c). The one at equilibrium, $f_{\mathrm{eq}}(t)$, is highlighted by a thick solid line. For $f(t)$ with finite $\dot{\gamma}$, all of them align with $f_{\mathrm{eq}}(t)$ at small $t$, and deviate from $f_{\mathrm{eq}}(t)$ at large $t$. We denote the point at which $f(t)$ deviates from $f_{\mathrm{eq}}(t)$ as $t_{\mathrm{c}}$, and mark them in figure 24 (c). Evidently, $t_{\mathrm{c}}$ can be viewed as the boundary between fast and slow groups: For $t<t_{\mathrm{c}}$, particles are not facilitated by external shear and should be classified as fast; while for $t>t_{\mathrm{c}}$, particles are facilitated and should be classified as slow. Then, we calculate the percentage of slow particles by $p_{\mathrm{slow}}^{\prime}=\int_{t_{\mathrm{c}}}^{\infty}f(t)\,\mathrm{d}t$, and show the result in figure 24 (d). The S-shaped dependence of $p_{\mathrm{slow}}^{\prime}$ on $\lg{\dot{\gamma}}$ is similar to the results of $p_{\mathrm{slow}}$ shown in figure 13, which are directly found from the MD results. The inset of figure 24 (d) shows $t_{\mathrm{c}}$ as a function of $\dot{\gamma}$ in the log-log scale, which reveals a clear relation of $t_{\mathrm{c}}\propto\dot{\gamma}^{-1}$. This behavior is also consistent with the MD results shown in figure 3 (c) and (d).
Figure 24: Functional forms of $f(t)$ as various $\dot{\gamma}$. (a) $f(t)$ in linear scale. (b) $f^{*}(t^{*})$, where $t^{*}=t/\bar{t}(E_{\mathrm{b}}^{\prime},\dot{\gamma})$ and $f^{*}=f/f_{\mathrm{max}}$. (c) $f(t)$ in log-log scale. The form at equilibrium $f_{\mathrm{eq}}(t)$ is highlighted by a thick line. For each $\dot{\gamma}$, we mark the position $t_{\mathrm{c}}$ at which $f(t)$ deviates from $f_{\mathrm{eq}}(t)$ by a circle with its color the same to that of the corresponding $\dot{\gamma}$. (d) The proportion of slow particles found from the model analysis $p_{\mathrm{slow}}^{\prime}$ as a function of $\dot{\gamma}$. Inset: $t_{\mathrm{c}}$ as a function of $\dot{\gamma}$ in log-log scale (symbols). The dashed line denotes the relation $t_{\mathrm{c}}\sim\dot{\gamma}^{-1}$.
@@ V.2 Steady shear
In the preceding subsection, we model $\tau_{\mathrm{LC}}$, $\tau_{\mathrm{sh}}$ and their relation as the interplay among local configuration, convection and thermal activation. The discussion is primarily based on the start-up deformation represented by the behavior with respect to a single basin. In this subsection, we will further leverage the activation picture to rationalize our findings on the steady shear.
@@
where $P=P(E_{\mathrm{b}},\gamma,t)$ is the probability that a coherent subsystem experiences a local energy barrier of $E_{\mathrm{b}}$ and a local strain of $\gamma$ at a given time $t$. In our picture, $P(E_{\mathrm{b}},\gamma,t)$ describes the dynamics at the particle level. The first term on the right-hand side of equation 19 represents the convective effect. The second term represents the activation from the distorted basin, where the energy barrier is given by $E_{\mathrm{b}}-E(\gamma)$. Here, the form of $E(\gamma)$ and the attempt frequency $\Gamma$ are set to those used in the preceding subsection. The third term represents the process that the yielded subsystems rearrange to new local equilibrium configurations. $\Gamma(t)$ is the total yielding rate, which is given by:
@@
$\rho(E_{\mathrm{b}})$ is the density of states for new barriers. Here, it is assumed that $\rho(E_{\mathrm{b}})$ does not depend on $\dot{\gamma}$ [23]. Therefore, $\rho(E_{\mathrm{b}})$ can be obtained from the dynamics of equilibrium state as follows. First, we measure the distribution of $\tau_{\mathrm{LC}}$ value at equilibrium $H_{\mathrm{eq}}(\tau_{\mathrm{LC}})$. Second, assuming the activation picture, we combine $H_{\mathrm{eq}}(\tau_{\mathrm{LC}})$ and equation 14 to find the distribution of barrier height $P_{\mathrm{b}}(E_{\mathrm{b}})$. Then, $\rho(E_{\mathrm{b}})$ can be found by:
@@
The model given by equations 1921 provides a framework for quantifying the flow behaviors under shear.
As demonstrated in section IV, the dynamics and rheology of supercooled liquids are strongly correlated to the local structure. It is interesting to understand this correlation from the shear-facilitated activation picture. We recall that local structural terms $g_{0,i}^{0}$ and $g_{2,i}^{-2}$ are most relevant to local dynamics compared to terms with other combinations of $l$ and $m$. In the present model, local dynamics is determined by the local energy barrier $E_{\mathrm{b}}$ and the local strain $\gamma$. As for $g_{0,i}^{0}$, it represents the local packing surrounding particle $i$. Notably, a recent work [28] has established the quantitative connection between the energy barrier for particles and the local packing efficiency characterized by the particle-wise radial particle density, which is similar to $g_{0,i}^{0}$, for the equilibrium state. Following this result, it is clear that the importance of $g_{0,i}^{0}$ for local dynamics comes from its remarkable relation to the local energy barrier $E_{\mathrm{b}}$. Moreover, we find that the terms with $l=0$ and $m=0$ alter with $\dot{\gamma}$ only slightly. This result might justify the above approximation that $\rho(E_{\mathrm{b}})$ is $\dot{\gamma}$-independent.
Our emphasis will be on local structural terms with $l=2$ and m=2m=-2. These terms have received less attention in previous studies of the equilibrium state. However, they are very important for shear states as shown in section IV. To understand this observation, we first refer to the conclusion that the strain of an affine shear $\gamma_{\mathrm{aff}}$ is proportional to $g_{2}^{-2}(r)$ [140]:
@@
where $g_{2}^{-2}(r)$ is the (l=2,m=2l=2,m=-2) coefficient of the spherical harmonic expansion of the global pair distribution function $g(\bm{r})$, and $g_{\mathrm{eq}}(r)$ is $g(\bm{r})$ at the undeformed state. Equation 22 suggests that $g_{2,i}^{-2}$ directly reflects the strain of the local configuration of particle $i$. Now, it is crucial to verify if the strain variable $\gamma$ in the activation picture, which can be calculated with the SGR model, can describe the configurational distortion reflected by $g_{2,i}^{-2}$ under steady shear. For this purpose, we perform the following calculations by employing the SGR model given by equations 1921.
Figure 25: (a) and (b) show the results of the structure feature $Y$, defined by equation 26, as a function of the distance $r$ at various $\dot{\gamma}$ for sample A and sample B, respectively. The first seven peaks of $Y(r)$ are marked. (c) and (d) show the magnitudes of the first seven peaks of $Y(r)$ (symbols) at different $\dot{\gamma}$ for sample A and sample B, respectively. The results of $\bar{\gamma}$ (dashed lines), defined by equation 24, at different $\dot{\gamma}$ are also plotted. In (c) and (d), the values of peak magnitudes of $Y(r)$ are scaled with a numerical factor to compare with $\bar{\gamma}$.
According to equation 19, it can be demonstrated that the distribution of local strain $\gamma$ under steady shear is given by the following equation [23, 139]:
@@
Here, we adopt the second generation of $g_{2,i}^{-2}$ to represent the local distortion, as suggested in section IV.1, and do not differentiate particle species when calculating $g_{2,i}^{-2}$ with equation 9. We construct the following vector for describing the distortion of the local configuration of particle $i$:
@@
We plot the results of $Y(r_{k})$ with a series of $k$ values at different $\dot{\gamma}$ for sample A and sample B in figure 25 (a) and (b), respectively. $Y(r_{k})$ displays an oscillation as $r_{k}$ increases, which is similar to the $r$-dependence of $g_{2}^{-2}(r)$ [44]. As expected, the amplitude of $Y$ grows with $\dot{\gamma}$. Figure 25 (c) and (d) display the magnitudes of the first seven peaks of $Y(r_{k})$, which are marked in figure 25 (a) and (b), as a function of $\dot{\gamma}$ for sample A and sample B, respectively. Meanwhile, we also plot the results of $\bar{\gamma}$ found from equation 24 for both samples in figure 25 (c) and (d). It is seen that the magnitudes of the peaks of $Y(r_{k})$ and $\bar{\gamma}$ exhibit highly consistent dependences on $\dot{\gamma}$. This consistency holds for different peaks of $Y(r_{k})$ within $r_{k}<3.5$. At larger $r_{k}$, this consistency weakens, implying the localization of such deformation coherency. The remarkable proportionality between $Y(r_{k})$ and $\bar{\gamma}$ within $r_{k}<3.5$ implies a nice correspondence between $g_{2,i}^{-2}$ and the local strain $\gamma$ in the present model.
In section IV.1, we show that local structural parameters constructed by $g_{0,i}^{0}$ and $g_{2,i}^{-2}$, which respectively reflect the energy barrier height $E_{\mathrm{b}}$ and the strain $\gamma$ of the particle-level basin according to the above discussion, well predict $\tau_{\mathrm{LC}}$ for all shear rates through a linear regression method. To explore the intrinsic mechanism of this correlation between local structure and dynamics, we will evaluate the dependence of $\tau_{\mathrm{LC}}$ on $\dot{\gamma}$ with the SGR model. Under steady shear, the distribution of local strain $\gamma$ and barrier height $E_{\mathrm{b}}$ in the SGR model is given by:
@@
Figure 26 gives $\bar{\tau}_{\mathrm{LC}}$ as a function of $\dot{\gamma}$ for both samples. We also plot the average $\bar{\tau}_{\mathrm{LC}}$ measured directly from MD data in figure 26. It is seen that the results found by the SGR model are well consistent with the MD results.
@@
Figure 26: (a) and (b) show $\bar{\tau}_{\text{LC}}$ (dashed lines), defined by equation 29, as a function of $\dot{\gamma}$ for sample A and sample B, respectively. Meanwhile, the average $\tau_{\mathrm{LC}}$ measured from the MD data (symbols) are also shown. For MD results, squares denote the results at the equilibrium state, and circles denote the results for finite shear rates.
@@
The results shown in figures 25 and 26 suggest a clear picture for the universal correlation between local structure and $\tau_{\mathrm{LC}}$ at all shear rates shown in section IV.1. The local structural terms related to $g_{0,i}^{0}$ and $g_{2,i}^{-2}$ respectively associate with the local energy barrier height $E_{\mathrm{b}}$ and the local strain $\gamma$ of the particle-level basin. $\tau_{\mathrm{LC}}$ is the waiting time to jump out from the basin characterized by $\gamma$ and $E_{\mathrm{b}}$. The convective effect is reflected by the tilt of the basin, as illustrated in figure 21. Under steady shear, the collective effects of thermal activation and convection are described by the SGR model. With these considerations, we can unify the values of $\tau_{\mathrm{LC}}$ from configurations at different shear rates.
@@
Figure 27: (a) and (b) show $P_{\mathrm{ac}}(\gamma_{\mathrm{ac}})$, given by equation 30, at various $\dot{\gamma}$ for sample A and sample B, respectively. Here, the values of $P_{\mathrm{ac}}$ are scaled as $P_{\mathrm{ac}}^{*}(\gamma_{\mathrm{ac}})=P_{\mathrm{ac}}(\gamma_{\mathrm{ac}})/P_{\mathrm{ac,max}}(\gamma_{\mathrm{ac}})$. The vertical dashed line in each panel denotes the effective range of $P_{\mathrm{ac}}(\gamma_{\mathrm{ac}})$ at the lowest $\dot{\gamma}$. (c) and (d) show the results of the percentage of shear-driven particles $p_{\mathrm{sd}}$ (dashed lines), defined by equation 31, at different $\dot{\gamma}$ for sample A and sample B, respectively. The results of the proportion of slow particles $p_{\mathrm{slow}}$, which has been shown in figure 13, are replotted (symbols).
@@
In our framework, the solid-liquid duality, which is based on the grouping of slow/fast particles, is the key to connecting $\tau_{\mathrm{LC}}$ to rheological behaviors. Especially, this grouping is also strongly correlated to the local structure featured by terms associated with $g_{0,i}^{0}$ and $g_{2,i}^{-2}$, as demonstrated in section IV.2. Thus, it is important to examine if the present model can give such grouping for conditions of steady shear. To explore this problem, we introduce the concept of activation strain $\gamma_{\mathrm{ac}}$, which represents the strain at which particles jump out from the basin. Particular attention is paid to the distribution of $\gamma_{\mathrm{ac}}$, denoted as $P_{\mathrm{ac}}(\gamma_{\mathrm{ac}})$. Under steady shear, $P_{\mathrm{ac}}(\gamma_{\mathrm{ac}})$ is given by:
@@
Figure 27 (a) and (b) give the $P_{\mathrm{ac}}(\gamma_{\mathrm{ac}})$ of sample A and sample B, respectively. For the sake of clarity, we plot the scaled results $P_{\mathrm{ac}}^{*}(\gamma_{\mathrm{ac}})=P_{\mathrm{ac}}(\gamma_{\mathrm{ac}})/P_{\mathrm{ac,max}}(\gamma_{\mathrm{ac}})$, where $P_{\mathrm{ac,max}}(\gamma_{\mathrm{ac}})$ is the maximum value of $P_{\mathrm{ac}}(\gamma_{\mathrm{ac}})$. At the lowest $\dot{\gamma}$ shown here (in the Newtonian regime), $P_{\mathrm{ac}}(\gamma_{\mathrm{ac}})$ is exponential-like. In figure 27 (a) and (b), we use vertical dashed lines to mark the effective range of $P_{\mathrm{ac}}(\gamma_{\mathrm{ac}})$ at the lowest $\dot{\gamma}$, denoted as $\gamma_{\mathrm{ac,c}}$. As $\dot{\gamma}$ increases, $P_{\mathrm{ac}}(\gamma_{\mathrm{ac}})$ shifts to the regime of larger $\gamma_{\mathrm{ac}}$. More importantly, $P_{\mathrm{ac}}(\gamma_{\mathrm{ac}})$ gradually exhibits a bimodal feature composed of an exponential-like component at small $\gamma_{\mathrm{ac}}$ regime and a gaussian-like component at large $\gamma_{\mathrm{ac}}$ regime, with the latter being enhanced by shear. The emergence of the gaussian-like component at large $\gamma_{\mathrm{ac}}$ manifests the change in the way of relaxation. For particles corresponding to the gaussian-like component, their local configurations are significantly strained by external shear before relaxation. Qualitatively, these particles are likely slow particles: Their dynamics are slower than external shear, so that there is enough time for their local configurations to be strongly strained by external shear. On the contrary, remaining particles can be considered as likely fast particles, since there is no enough time for external shear to drive their local configurations to larger strains.
As seen in figure 27 (a) and (b), $\gamma_{\mathrm{ac,c}}$ gives a rough boundary between the exponential-like component and the gaussian-like component. Thus, we can evaluate the percentage of the gaussian-like component as follows: